OpenStudy (ellableuuu):

Find f(6) from the table below. y f(y) 3 5 4 6 5 4 6 3 - Select one: a. 3 b. 4 c. 5 d. 6

OpenStudy (welshfella):

f(6) means the value of the function when y = 6

OpenStudy (welshfella):

so read it off the table

OpenStudy (ellableuuu):

so 4?

OpenStudy (welshfella):

No f(5) = 4

OpenStudy (ellableuuu):

I'm confused

OpenStudy (welshfella):

what is the value of f(y) when y = 6?

OpenStudy (welshfella):

3

OpenStudy (welshfella):

- check the table

OpenStudy (ellableuuu):

is it 3 because you multiply it to get 6?

OpenStudy (welshfella):

NO its 3 because it is given as 3 in the table.

OpenStudy (ellableuuu):

ohhhh sorry, i get it

OpenStudy (ellableuuu):

i know what i was doing wrong, i kept looking at the 6 in the f(y) column

OpenStudy (ellableuuu):

i wasn't thinking oops

OpenStudy (welshfella):

lets take a simple function f(y) = 2y when y = 2 , f(y) = 2*2 = 4 and we say that f(2) = 4 In the same way f(3) = 2*3 = 6 get it?

OpenStudy (ellableuuu):

I get it, thank you.

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

its just the notation.